A Complete Bibliography of the Journal of Time Series Econometrics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 August 2024 Version 1.07

Title word cross-reference

AR(1) [Sim12]. b [ASV09]. D [TZCP19]. I(1) [Eve10]. Z^{d} [DF22].

-Vine [TZCP19].

across [Mar21]. Activations [LXZ13]. Adapted [OLL19, PMA12]. Adaptive [LX10]. Additive [HMS11]. Adequacy [WSAF20]. Adjustment [BD15, Can22, Oll21]. Aggregate [Qui21]. Aggregates [HT13]. Algebra [Sim12]. Algorithm [Koc11]. Analysing [AS18]. Analysis [BiS09, BPS21, CP22, KMSZ19, Lee19, Man24, Pol09]. Analyzing [KSJ17]. Annual [LX11]. Antipersistent [Sim12]. APARCH [JP17]. Application [ABH18, APMA20, BiS09, GJ23, PCG12, Vaf15]. Applications [OLL19, TZCP19]. Applied [GW13]. Approach [BFFB19, DL23, KSJ17, KP21, OLL19, Qui21, Sim12]. Approximations [Man10]. Approximators [Koc11]. ARCH [AL16, KR09, MY13].

ARFIMA [Bou18, Man10]. ARMA [DF22]. ARMA-type [DF22]. Artificial [LXZ13]. Asset [TTS22]. Asymmetry [AS15]. Asymptotic [HT13, HMX13, TK23]. Asymptotically [Oku14]. Asymptotics [KR09]. Augmented [AHL17]. Australia [Kol19]. Autocorrelated [ASV09, PMA12, SKT17, Poi18]. Autocorrelations [Oku14, WH11]. Autocovariances [Oku14]. Automatic [CDH11]. Autoregression [AUY20, Cai09]. Autoregressions [AS18, Akn13, LS11]. Autoregressive [Dem09, LEP12, Ngu16]. Average [BZ14, ML12, Ngu16, SKJ19]. Averaging [DL23, GZ20].

Based [BD16, DR15, GW13, LL10, LX10, Mil10]. Basis [GW13]. Bayesian [AC20]. Be [DI11]. Behavior [HT13, JP17, KP21]. BEKK [AM22, AS23, Bur15, HLP16]. BEKK-CAW [AS23]. benchmark [Vaf15]. Better [TKI21]. between [Ngu16]. Bias [BZ14, DK19, LEP12, Skr14]. Biases [AL12]. BINMA [KSJ17]. Binomial [KSJ18]. Bivariate [KSJ18, SKJ19]. Block [PPP15]. Bond [GW13]. Boom [CP22]. Bootstrap [LEP12, PPP15, Wan14]. Brazilian [CP22]. Break [Skr14, Sol16]. Breaks [BiS09, BFC11, Skr18, Smi12, TK23]. Business [Kol19].

capital [Sin16]. Carlo [Bur15, Non16]. Case [PR11]. CAW [AS23]. CDS [GW13]. CDS-Bond [GW13]. chain [Non16]. Change [ILT17, PCG12, TK23]. Changed [GP10]. Changes [Kur15]. Changing [HMX13]. Checking [WSAF20]. Chow [Poi18]. Chow-Lin [Poi18]. Cities [BiS09]. Clarifications [Pol09]. Class [AD14]. Cloning [Lau13]. Close [Lar15]. Co [OG21]. Co-integrated [OG21]. Coefficient [Akn13]. Cointegrated [APMA20]. Cointegrating [Can22]. Cointegration [BD15, GW13, HI12, ML12]. COLS [LEP12]. Commodities [CP22]. Commodity [CP22, Man24]. Common [CHP11, HLP16]. Comparison [Sha20]. component [Non16]. Components [BD15, CHP11, MM14]. Computing [ABH18]. Condition [AHL17, Skr18]. Conditional [BZ14, San17]. Consideration [WG11]. Constrained [Bur15]. Consumption [Qui21]. Continuous [OG21, Sim12]. Convergence [BiS09, JL10]. Copula [TZCP19]. Copulas [MTCdM11]. corrected [SKT17]. Correction [DK19, Skr14]. Correlated [MM14]. Correlation [AS15]. Correlograms [AL12]. Costationarity [CN10]. Count [WSAF20]. Covariances [San17]. Covariate [AHL17, GW13]. Critical [ASV09]. Cumulative [KMSZ19]. Curve [Sha20]. Curves [KMSZ19]. Cycles [Kol19, Pol14].

Daily [CGW11, Oll21]. Data [BD16, BiS09, Lau13, LX11, NB09, Oku14]. Data-Driven [BD16]. Dependence [JP17, LL10]. Dependent [AD14, MA22a, MA22b, Sha20]. Design [PMA12]. Design-Adapted [PMA12]. Detecting [CHP11]. Detection [HMS11]. Deterministic [BD15, VSGZ10, WH11]. Dickey [Sol16]. Dickey-Fuller [Sol16]. Different [DF22]. Directional [MR22]. Discrete [OG21]. Discretization [Sim12]. Dispersed [SKJ19]. dispersion [KSJ18]. Distortion [Skr14]. Distributed [AK18]. Distribution [TK23]. Distributions [SN09]. Disturbances [Ngu16]. Diverse [TM17]. Divorce [KP21]. Do [Lip17]. Domain [HT13, HPW21]. Double [LL10]. Double-Length [LL10]. Drift [Lar15]. Driven [BD16, KSJ18, Sim12]. Dynamic [AS15, DL23, TZCP19, Poi18]. Dynamics [APMA20, CHP11].

ECCC [Xu24]. ECCC-GARCH [Xu24]. Econometric [HM11]. Economic [BCHL11, LS11]. Economy [CP22]. Edge [DF22]. Edge-Effect [DF22]. Edgeworth [AD14]. Editors [BCHL11]. Effect [DF22, JX18, PCG12]. Effects [MY13, Oku14]. Efficiency [AS18]. Efficient [JN11]. EGARCH [DK19]. either [Sim12]. Elections [Kol19]. Electricity [Lip17]. ELG [CP22]. Elliptical [SN09]. Ensemble [TKI21]. Environment [NB09]. Equity [AS18]. Error [LEP12, MA22a, MA22b, Xu24]. Errors [ABH18, JX18, ML12, Man24, PMA12, SKT17]. Estimating [MA22a, MR22, TK23, WH11]. Estimation [Akn13, ABH18, AC20, BZ14, CP20, Eve10, HI12, HPW21, LX10, LEP12, LE23, MTCdM11, Oku14, OG21, OLL19, Xu24]. Estimator [Lau13, RW10, TK23]. Estimators [DK19, Sha20]. Euro [PCG12]. Evaluating [CDH11, JS11]. Evidence [VSGZ10, Sin16]. Exact [Sim12]. Examining [Pol14]. Exchange [TKI21]. Expansions [AD14]. Expectations [JS11]. Expected [Qui21]. Exponent [Sha20]. Exponential [AS15]. Export [CP22]. Export-Led [CP22]. Expression [BFFB19]. Extended [Man10, Poi18]. Extraction [MW10, MM14, TM17].

Factor [Kaw23]. Factors [HLP16]. Fama [Vaf15]. Fama-French [Vaf15]. FCVAR [Qui21]. Feasible [BZ14]. Fill [TK23]. Filters [WM16]. Financial [BCHL11]. Finite [DK19]. Finite-Sample [DK19]. First [BZ14, HI12]. First-Order [BZ14]. Fiscal [Man24]. Fit [MA22b, MW10]. Fixed [ASV09, Sol16]. Fixed- [ASV09]. Flexible [AUY20, KSJ18]. Flow [OG21]. Follow [AS18]. Forecasting [BFFB19, BCHL11, DL23, Koc11, LX11, OLL19, SN09, TKI21, Vaf15]. Forecasts [BZ14, GZ20]. Fossil [Lip17]. Fourier [Pol09]. Fractional [BD16, Bou18, HI12, ILT17, Lar15, Man10, Sim12]. French [Vaf15]. Frequency [AS18, AUY20, HT13, HPW21]. Frontmatter [Ano14a, Ano14b, Ano15a, Ano15b, Ano16a, Ano16b, Ano21a, Ano21b, Ano22a, Ano22b, Ano23a, Ano23b, Ano24]. Fuels [Lip17]. Full [KSJ17]. Fuller [Sol16]. Functional [AC20]. Functions [MR22].

GARCH [AM22, Bur15, CGW11, CP20, JL10, Xu24]. GARCH-BEKK [AM22]. Gaussian [Cai09, Man10, Man24]. Gegenbauer [HPW21]. Gene [BFFB19]. General [BD15, HPW21]. Generalized [APMA20, Bou18, LBN20, OLL19]. Generation [Lip17]. Global [BPS21].

Good [RW10]. Goodness [MA22b, MW10]. Goodness-of-Fit [MA22b, MW10]. GQL [KSJ17]. Graphical [JS11]. Growth [CP22].

Haar [PMA12]. Hamiltonian [Bur15]. Heterogeneous [PCG12]. Hidden [LXZ13]. Highly [ASV09]. Homogeneous [Kur15, KSJ18]. House [Mar21]. Hurst [Sha20]. Hybrid [Lau13, CGW11]. Hyper [AM22]. Hyper-Rotated [AM22]. Hypotheses [Can22]. Hypothesis [CP22, DR15].

Idea [Pol14]. Identifying [MY13]. Ideology [Kol19]. Impact [AHL17, Lip17]. Implications [TKI21]. Impossibility [PR11]. Improved [BD15, Ngu16]. Improving [LE23]. Impulse [MR22]. Impulse-Response [MR22]. In-Fill [TK23]. INAR [KSJ18, LBN20]. Increased [Lip17]. Increment [BD16]. Index [LX11]. Individual [Oku14]. Indonesian [Man24]. Inefficient [AS18]. Inference [DF22, Eve10]. Infinite [Dem09]. Infinite-Order [Dem09]. Inflated [LBN20]. Inflated-parameter [LBN20]. Inflation [DL23, GP10, LX11]. Initial [AHL17, Skr18]. Innovations [Cai09, LBN20, Sim12]. Innovative [OLL19]. Instrumental [NB09]. Instruments [Dem09]. integrated [OG21]. Integration [Bou18, ILT17, dO11]. Interest [APMA20]. International [Sin16]. Interpretations [Pol09]. Intervals [CP20]. Intra [CGW11]. Intra-Daily [CGW11]. Intraday [GHK13, KMSZ19]. Introduction [BCHL11]. Irrelevance [PR11]. IV [Mil10].

Jackknife [LEP12]. Just [RW10].

KPSS [ASV09, Skr14].

Lag [AC20, LL10]. Lagged [OLL19]. Laplace [JX18]. Learning [Koc11]. Least [Akn13]. Led [CP22]. Length [LL10]. Level [BiS09]. Levels [KSJ18, LX11]. Leverage [JX18]. Light [Lip17]. Likelihood [DK19, GJ23, JN11, Lau13, Mil10, Xu24]. Likelihood-Based [Mil10]. Limit [AL16]. Lin [Poi18]. Linear [KR09, WM16]. Linkages [Mar21]. Local [GJ23, OLL19]. Locally [AD14, CN10]. Logs [LX11]. Long [AS15, APMA20, DR15, Mil10, PR11, Sha20, Sim12]. Long-range [Sha20]. Long-run [PR11].

Macroeconomic [Mar21, MR22]. Many [LXZ13]. Market [AS18, Vaf15]. Markov [Non16, Smi12]. Masthead [Ano13a, Ano13b]. Matrix [Aba12, AS15]. Matrix-Exponential [AS15]. Matter [Lip17]. Maximum [DK19, GJ23, Lau13, Xu24]. May [DI11]. Mean [ILT17, LEP12]. Measure [MW10]. Memory [AS15, APMA20, DR15, Sim12]. Meta [SN09]. Meta-Elliptical [SN09]. Method [AS18, HPW21, Lee19, OLL19, Poi18]. Methods [ABH18]. Misspecified [JL10]. Mixed [AUY20]. Mixed-Frequency [AUY20]. mobility [Sin16]. Model [AL16, AC20, BZ14, Bou18, CDH11, DK19, DL23, GJ23, JP17, Sim12, TZCP19, WSAF20, Xu24]. Modeling [DI11, GS14, JX18, Mar21, TKI21]. Modelling [HM11, SKJ19]. Models [AS23, BPS21, CGW11, HLP16, JL10, JS11, Kaw23, KS11, KR09, Lau13, LEP12, LE23, MA22a, MA22b, MY13, MR22, Ngu16, Smi12, Non16, Poi18]. Moments [OLL19]. Mongolia [DL23]. Monitoring [GHK13, WH13]. Monotonicity [TTS22]. Monte [Bur15, Non16]. Monthly [LX11]. Moving [BZ14, ML12, Ngu16, SKJ19]. Multidimensional [BD16]. Multiple [BiS09, Kur15]. Multiplicative [Xu24]. Multivariate [AM22, Mil10].

Nearly [JN11]. Negative [KSJ18]. Neglected [LXZ13]. Network [Lee19]. Networks [LXZ13]. Neural [LXZ13, Lee19]. Noise [Man10]. Non [AL16, BCHL11, Cai09, KSJ18, KR09, Kur15, Man24, GZ20]. Non-Gaussian [Cai09, Man24]. Non-Homogeneous [Kur15, KSJ18]. Non-Linear [KR09]. Non-Parametric [GZ20]. Non-stationarity [BCHL11]. Non-stationary [AL16]. Noncausal [LS11]. Nonlinear [Dem09, Lee19, Mil10]. Nonlinearity [LXZ13, LL10, San17]. Nonparametric [BFC11, dO11]. Nonstationary [Akn13, DI11, TM17]. Note [AL16]. NoVaS [CP20]. null [Sol16]. Numerical [ABH18].

Observation [KSJ18]. Observation-Driven [KSJ18]. Observations [HM11]. Off [DI11]. Oil [BFFB19]. Omitted [Eve10]. One [TK23]. Optimal [MM14, Wan14, WM16]. Order [BZ14, Dem09]. Outliers [HMS11]. Outlying [HM11]. Over-dispersion [KSJ18]. Ox [Non16].

Panel [BiS09, Dem09, Oku14]. Panels [Mil10, PCG12]. Parameter [Bou18, LBN20]. Parameters [HMX13]. Parametric [GZ20]. Parity [BiS09]. Partially [LX10]. Particle [Non16]. Particular [AS18]. Party [Kol19]. Pattern [GHK13]. PCSE [RW10]. Pearson [WSAF20]. Penalized [GZ20]. Period [CP22]. Periodic [dO11]. Periodicity [BCHL11, CGW11]. Point [CP20, TK23, Wan14]. Policy [Man24, TKI21]. Political [Kol19]. Power [ASV09, BiS09, LBN20]. Prediction [CP20, WM16]. Predictions [LE23]. Prequential [SN09]. Presence [ML12, Oku14, WH11, Sol16]. Price [BFFB19, BiS09, LX11, Man24, Mar21]. Prices [AS18, Lip17]. Prior [AUY20]. Problems [WM16]. Process [APMA20, KSJ17, Lar15]. Processes [AD14, AS15, Dem09, DF22, HPW21, LBN20, SKJ19]. Programming [BFFB19]. Proportion [AS18]. Purchasing [BiS09].

QMLE [AL16, JL10, KR09]. Quantile [GZ20]. Quantiles [MR22]. Quasi [Xu24].

Random [Akn13, Lar15]. Randomized [LXZ13]. range [Sha20]. Rank [BD15, Mil10]. Rate [TKI21]. Rates [APMA20, KP21]. Ratio [BD16, JN11]. Rational [JS11]. Re [GS14]. Re-switching [GS14]. Real [WH13, WM16].

Real-Time [WH13, WM16]. Realized [AS23, Kaw23, WH13]. Recursive [BD15, Sol16]. Reducing [Skr14]. Regression [BPS21, KP21, Ngu16, PMA12, Smi12]. Regressions [HMX13, LL10, SKT17]. Regressors [Kur15]. Renewable [Lip17]. Representation [HLP16, Xu24]. Representations [AL12]. Residual [GW13]. Residual-Based [GW13]. Residuals [WSAF20, Poi18]. Response [MR22]. Results [KS11]. Return [CGW11, DI11, KMSZ19, Vaf15]. Returns [BFFB19, Qui21, TTS22]. Review [ABH18]. Revision [MW10]. Rich [NB09]. right [Sol16]. right-tailed [Sol16]. Risk [ABH18, KMSZ19, TZCP19, Vaf15]. risk-return [Vaf15]. Robust [KSJ17, TTS22, Sin16]. Root [Aba12, AHL17, BFC11, Dem09, LX10, PPP15, Skr18, VSGZ10, Wan14]. Roots [BD16, JN11]. Rotated [AM22]. Rotation [GS14]. Rules [TM17]. run [PR11].

Sample [Can22, DK19]. Sampling [TM17]. Seasonal [HMS11, JN11, LX11, Oll21]. Seemingly [SKT17]. Selecting [NB09]. Selection [BPS21, CDH11, Ngu16]. Self [DR15]. Self-Similarity [DR15]. Semantics [Pol14]. Semiparametric [BD16]. Sensitivity [BPS21]. Sequences [AD14]. Sequential [AK18]. Series [AL12, ASV09, BCHL11, CN10, Eve10, HMS11, HM11, KSJ17, LS11, Lee19, LE23, LBN20, Mil10, MTCdM11, Oll21, Sha20, TM17, WSAF20, WG11, Non16, Sin16]. Signal [MW10, MM14, TM17]. Significance [SKT17]. Similarity [DR15]. Simple [Kaw23, Man10]. SIS [GJ23]. Size [ASV09, AK18, Skr14, SKT17]. Small [Can22, LE23]. Smooth [AD14, Bou18, KP21]. Smoothly [HMX13]. Some [JS11, KS11]. SPARMA [MA22a, MA22b]. Spatial [AC20, LL10]. Spurious [Pol14]. Square [Aba12]. Squared [LEP12]. Squares [Akn13]. Stage [Akn13, HI12]. Standard [ABH18]. State [AUY20]. States [Mar21, Sin16]. Stationarity [BD16, BCHL11]. Stationary [AL12, CN10, KSJ18, AL16]. Statistical [Pol09]. Statistics [BD16]. Steady [AUY20]. Steady-State [AUY20]. Still [Lip17]. Stochastic [Kaw23, Lau13]. Stock [OG21, Qui21]. Structural [BiS09, BFC11, Kur15, MY13, PCG12, Skr14]. Structure [JP17]. Style [GS14]. SVAR [Man24]. Switching [GS14]. Syllogisms [Pol14]. Synchronization [Mar21]. Systems [OG21].

Tail [JP17]. tailed [Sol16]. Tapered [PPP15]. targeting [Bur15]. Technique [JS11]. techniques [Non16]. Temporal [HT13]. Temporally [GJ23]. Terms [MA22a, MA22b]. Test [ASV09, BFC11, DR15, GW13, Mil10, Ngu16, Skr14, TTS22, WH13]. Testing [AK18, Can22, Dem09, ILT17, Kur15, LXZ13, LX10, ML12, PPP15, PCG12, San17, Skr18, VSGZ10]. Tests [AHL17, BD16, BD15, JN11, LL10, MA22b, SKT17, Wan14, dO11, Sin16, Sol16]. their [Mar21]. Theorems [PR11]. Theory [AL16, DK19, HMX13]. There [VSGZ10]. Think [RW10]. Threshold [AS18, KS11]. Time [AS18, ASV09, BCHL11, CN10, CP20, Eve10, HMS11, HM11, KSJ17, LS11, Lee19, LE23, Mar21, Mil10, MTCdM11, Oku14, OG21, Oll21, SN09, Sha20, Sim12, TK23, TM17, WH13, WSAF20, WG11, WM16, Non16, Sin16]. time-series [Sin16]. Time-Varying [Mar21, CP20]. Trade [DI11, PCG12]. Trade-Off [DI11]. tradeoff [Vaf15]. Transformations [AD14]. Transition [Bou18, KP21]. Transitory [CHP11]. Trend [Skr18, VSGZ10]. Trends [WH11, WG11]. Two [Akn13]. Two-Stage [Akn13]. type [DF22].

U.S. [BiS09, GP10, Mar21]. Unbiased [Oku14]. Underlying [LX11]. Understanding [DF22]. Uniform [AD14]. Uniformly [AK18]. Unit [AHL17, BD16, BFC11, Dem09, JN11, LXZ13, LX10, PPP15, Skr18, VSGZ10, Wan14]. United [Sin16]. Univariate [HLP16]. Universal [Koc11]. unobserved [Non16]. Unrelated [SKT17]. Using [ASV09, BFFB19, BPS21, KSJ17, LXZ13, LX11, PMA12, TKI21, WSAF20, MR22, Non16, Xu24].

Valid [AD14]. Validation [SN09]. Validity [CP22]. Value [ABH18, TZCP19]. Value-at-Risk [ABH18, TZCP19]. Values [ASV09]. VaR [TZCP19, MY13]. Variable [BPS21]. Variables [Eve10, NB09, OG21]. Variance [PR11]. Variances [MW10]. Varying [Mar21, SN09, CP20]. Vector [AUY20, Xu24]. Vectors [Can22]. Versus [CP20, LX11]. via [MY13]. Vine [TZCP19]. Volatility [BFFB19, CP20, DI11, GHK13, GP10, JX18, Kaw23, Lau13, WH13, Vaf15].

Walk [Lar15]. Wavelet [MTCdM11]. Wavelets [PMA12]. way [DF22]. Weakly [AD14]. Wealth [Qui21]. Weighted [Akn13].

References

Abadir:2012:SRM

[Aba12] Karim M. Abadir. The square root of a matrix. Journal of Time Series Econometrics, 4(2):4:1–4:5, November 2012. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

           URL http://www.degruyter.com/view/j/jtse.2012.4.issue-
           2/1941-1928.1140/1941-1928.1140.xml.

Ardia:2018:MCN

[ABH18] David Ardia, Keven Bluteau, and Lennart F. Hoogerheide. Methods for computing numerical standard errors: Review and application to value-at-risk estimation. Journal of Time Series Econometrics, 10(2):20170011:1–20170011:9, July 2018. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

           tronic). URL http://www.degruyter.com/view/j/jtse.2018.
           10.issue-2/jtse-2017-0011/jtse-2017-0011.xml.

Aw:2020:BEF

[AC20] Alassane Aw and Emmanuel Nicolas Cabral. Bayesian estimation of the functional spatial lag model. Journal of Time Series Econometrics, 12(2):??, ???? 2020. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/jtse-2019-0047/html.

Arvanitis:2014:VLU

[AD14] Stelios Arvanitis and Antonis Demos. Valid locally uniform Edgeworth expansions for a class of weakly dependent processes or sequences of smooth transformations. Journal of Time Series Econometrics, 6(2):183–235, July 2014. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2014.6.issue-
2/jtse-2012-0003/jtse-2012-0003.xml.

Aristidou:2017:IIC

[AHL17] Chrystalleni Aristidou, David I. Harvey, and Stephen J. Leybourne. The impact of the initial condition on covariate augmented unit root tests. Journal of Time Series Econometrics, 9(1):20150013:1–20150013:23, January 2017. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2017.9.issue-
1/jtse-2015-0013/jtse-2015-0013.xml.

Anatolyev:2018:STU

[AK18] Stanislav Anatolyev and Grigory Kosenok. Sequential testing with uniformly distributed size. Journal of Time Series Econometrics, 10(2):20170002:1–20170002:22, July 2018. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2018.
10.issue-2/jtse-2017-0002/jtse-2017-0002.xml.

Aknouche:2013:TSW

[Akn13] Abdelhakim Aknouche. Two-stage weighted least squares estimation of nonstationary random coefficient autoregressions. Journal of Time Series Econometrics, 5(1):25–46, May 2013. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2013.5.issue-
1/jtse-2012-0011/jtse-2012-0011.xml.

Abadir:2012:BCA

[AL12] Karim M. Abadir and Rolf Larsson. Biases of correlograms and of AR representations of stationary series. Journal of Time Series Econometrics, 4(1):1:1–1:11, May 2012. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2012.4.issue-
1/1941-1928.1130/1941-1928.1130.xml.

Arvanitis:2016:NQL

[AL16] Stelios Arvanitis and Alexandros Louka. A note on the QMLE limit theory in the non-stationary ARCH(1) model. Journal of Time Series Econometrics, 8(1):21–39, January 2016. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2016.8.issue-
1/jtse-2014-0034/jtse-2014-0034.xml.

Asai:2022:MHR

[AM22] Manabu Asai and Michael McAleer. Multivariate hyper-rotated GARCH-BEKK. Journal of Time Series Econometrics, 14(2):175– 198, July 2022. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/jtse-2021-0006/html.

Anonymous:2013:Ma

[Ano13a] Anonymous. Masthead. Journal of Time Series Econometrics, 5 (1):i, May 2013. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL http://www.degruyter.com/view/j/jtse.
2013.5.issue-1/jtse-2013-masthead1/jtse-2013-masthead1.
xml.

Anonymous:2013:Mb

[Ano13b] Anonymous. Masthead. Journal of Time Series Econometrics, 5 (2):i, November 2013. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL http://www.degruyter.com/view/j/
jtse.2013.5.issue-2/jtse-2013-masthead2/jtse-2013-masthead2.
xml.

Anonymous:2014:Fa

[Ano14a] Anonymous. Frontmatter. Journal of Time Series Econometrics, 6(1):i, January 2014. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL http://www.degruyter.com/view/j/
jtse.2014.6.issue-1/jtse-2014-masthead1/jtse-2014-masthead1.
xml.

Anonymous:2014:Fb

[Ano14b] Anonymous. Frontmatter. Journal of Time Series Econometrics, 6(2):i, July 2014. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2014.6.issue-2/jtse-2014-frontmatter2/
jtse-2014-frontmatter2.xml.

Anonymous:2015:Fa

[Ano15a] Anonymous. Frontmatter. Journal of Time Series Econometrics, 7(1):i, January 2015. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2015.7.issue-1/jtse-2015-frontmatter1/
jtse-2015-frontmatter1.xml.

Anonymous:2015:Fb

[Ano15b] Anonymous. Frontmatter. Journal of Time Series Econometrics, 7(2):i, July 2015. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2015.7.issue-2/jtse-2015-frontmatter2/
jtse-2015-frontmatter2.xml.

Anonymous:2016:Fa

[Ano16a] Anonymous. Frontmatter. Journal of Time Series Econometrics, 8(1):i, January 2016. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2016.8.issue-1/jtse-2016-frontmatter1/
jtse-2016-frontmatter1.xml.

Anonymous:2016:Fb

[Ano16b] Anonymous. Frontmatter. Journal of Time Series Econometrics, 8(2):i, July 2016. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2016.8.issue-2/jtse-2016-frontmatter2/
jtse-2016-frontmatter2.xml.

Anonymous:2021:Fa

[Ano21a] Anonymous. Frontmatter. Journal of Time Series Econometrics, 13(1):i–iii, January 22, 2021. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/jtse-2021-frontmatter1/html.

Anonymous:2021:Fb

[Ano21b] Anonymous. Frontmatter. Journal of Time Series Econometrics, 13(2):i–ii, July 2021. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL https://www.degruyter.com/document/
doi/10.1515/jtse-2021-frontmatter2/html.

Anonymous:2022:Fa

[Ano22a] Anonymous. Frontmatter. Journal of Time Series Econometrics, 14(1):i–ii, January 2022. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/jtse-2022-frontmatter1/html.

Anonymous:2022:Fb

[Ano22b] Anonymous. Frontmatter. Journal of Time Series Econometrics, 14(2):i–ii, July 2022. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL https://www.degruyter.com/document/
doi/10.1515/jtse-2022-frontmatter2/html.

Anonymous:2023:Fa

[Ano23a] Anonymous. Frontmatter. Journal of Time Series Econometrics, 15(1):i–ii, January 2023. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/jtse-2023-frontmatter1/html.

Anonymous:2023:Fb

[Ano23b] Anonymous. Frontmatter. Journal of Time Series Econometrics, 15(2):i–ii, July 2023. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL https://www.degruyter.com/document/
doi/10.1515/jtse-2023-frontmatter2/html.

Anonymous:2024:F

[Ano24] Anonymous. Frontmatter. Journal of Time Series Econometrics, 16(1):i–ii, January 2024. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/jtse-2024-frontmatter1/html.

Asai:2020:CDG

[APMA20] Manabu Asai, Shelton Peiris, Michael McAleer, and David E. Allen. Cointegrated dynamics for a generalized long memory process: Application to interest rates. Journal of Time Series Econometrics, 12(1):??, January 2020. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://www.degruyter.

com/view/journals/jtse/12/1/article-20180024.xml.

Asai:2015:LMA

[AS15] Manabu Asai and Mike K. P. So. Long memory and asymmetry for matrix-exponential dynamic correlation processes. Journal of Time Series Econometrics, 7(1):69–94, January 2015. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2015.7.issue-
1/jtse-2013-0012/jtse-2013-0012.xml.

Ahmed:2018:WPT

[AS18] Muhammad Farid Ahmed and Stephen Satchell. What proportion of time is a particular market inefficient? … A method for analysing the frequency of market efficiency when equity prices follow threshold autoregressions. Journal of Time Series Econometrics, 10(2):20160021:1–20160021:22, July 2018. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2018.
10.issue-2/jtse-2016-0021/jtse-2016-0021.xml.

Asai:2023:RBC

[AS23] Manabu Asai and Mike K. P. So. Realized BEKK-CAW models. Journal of Time Series Econometrics, 15(1):49–77, January 2023. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2022-0009/html.

Amsler:2009:KTU

[ASV09] Christine Amsler, Peter Schmidt, and Timothy J. Vogelsang. The KPSS test using fixed-b critical values: Size and power in highly autocorrelated time series. Journal of Time Series Econometrics, 1(1):??, January 2009. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2009.1.1/jtse.2009.1.1.1027/jtse.2009.1.1.1027.
xml.

Ankargren:2020:FMF

[AUY20] Sebastian Ankargren, Måns Unosson, and Yukai Yang. A flexible mixed-frequency vector autoregression with a steady-state prior. Journal of Time Series Econometrics, 12(2):??, ???? 2020. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2018-0034/html.

Bollerslev:2011:PNS

[BCHL11] Tim Bollerslev, Bent Jesper Christensen, Niels Haldrup, and Asger Lunde. Periodicity, non-stationarity, and forecasting of economic and financial time series: Editors’ introduction. Journal of Time Series Econometrics, 3(1):1:1–1:8, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2011.
3.1/jtse.2011.3.1.1098/jtse.2011.3.1.1098.xml.

Born:2015:RAG

[BD15] Benjamin Born and Matei Demetrescu. Recursive adjustment for general deterministic components and improved cointegration rank tests. Journal of Time Series Econometrics, 7(2):143–179, July 2015. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2015.
7.issue-2/jtse-2013-0005/jtse-2013-0005.xml.

Bardet:2016:SSF

[BD16] Jean-Marc Bardet and Béchir Dola. Semiparametric stationarity and fractional unit roots tests based on data-driven multidimensional increment ratio statistics. Journal of Time Series Econometrics, 8(2):115–153, July 2016. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2016.8.issue-2/jtse-2014-0031/jtse-2014-
0031.xml.

Belaire-Franch:2011:NUR

[BFC11] Jorge Belaire-Franch and Dulce Contreras. Nonparametric unit root test and structural breaks. Journal of Time Series Econometrics, 3(2):3:1–3:14, April 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2011.3.2/jtse.2011.3.2.1048/jtse.2011.
3.2.1048.xml.

Baffour:2019:FVR

[BFFB19] Alexander Amo Baffour, Jingchun Feng, Liwei Fan, and Beryl Adormaa Buanya. Forecasting volatility returns of oil price using gene expression programming approach. Journal of Time Series Econometrics, 11(2):??, July 2019. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL

https://www.degruyter.com/view/j/jtse.2019.11.issue-2/
jtse-2017-0022/jtse-2017-0022.xml.

Basher:2009:PLC

[BiS09] Syed A. Basher and Josep Lluís Carrion i Silvestre. Price level convergence, purchasing power parity and multiple structural breaks in panel data analysis: An application to U.S. cities. Journal of Time Series Econometrics, 1(1):??, January 2009. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2009.
1.1/jtse.2009.1.1.1000/jtse.2009.1.1.1000.xml.

Boubaker:2018:GAM

[Bou18] Heni Boubaker. A generalized ARFIMA model with smooth transition fractional integration parameter. Journal of Time Series Econometrics, 10(1):20150001:1–20150001:21, January 2018. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2018.
10.issue-1/jtse-2015-0001/jtse-2015-0001.xml.

Becker:2021:VSR

[BPS21] William Becker, Paolo Paruolo, and Andrea Saltelli. Variable selection in regression models using global sensitivity analysis. Journal of Time Series Econometrics, 13(2):187–233, July 2021. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2018-0025/html.

Burda:2015:CHM

[Bur15] Martin Burda. Constrained Hamiltonian Monte Carlo in BEKK GARCH with targeting. Journal of Time Series Econometrics, 7 (1):95–113, January 2015. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2015.7.issue-1/jtse-2013-0013/jtse-2013-0013.xml.

Bao:2014:EBF

[BZ14] Yong Bao and Ru Zhang. Estimation bias and feasible conditional forecasts from the first-order moving average model. Journal of Time Series Econometrics, 6(1):63–80, January 2014. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2014.6.issue-
1/jtse-2013-0015/jtse-2013-0015.xml.

Cai:2009:ANG

[Cai09] Yuzhi Cai. Autoregression with non-Gaussian innovations. Journal of Time Series Econometrics, 1(2):??, December 2009.

CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2009.
1.2/jtse.2009.1.2.1016/jtse.2009.1.2.1016.xml.

Canepa:2022:SSA

[Can22] Alessandra Canepa. Small sample adjustment for hypotheses testing on cointegrating vectors. Journal of Time Series Econometrics, 14(1):51–85, January 2022. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/jtse-2020-0044/html.

Castle:2011:EAM

[CDH11] Jennifer L. Castle, Jurgen A. Doornik, and David F. Hendry. Evaluating automatic model selection. Journal of Time Series Econometrics, 3(1):8:1–8:33, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.

degruyter.com/view/j/jtse.2011.3.1/jtse.2011.3.1.1097/
jtse.2011.3.1.1097.xml.

Chen:2011:HGM

[CGW11] Xilong Chen, Eric Ghysels, and Fangfang Wang. HYBRID GARCH models and intra-daily return periodicity. Journal of Time Series Econometrics, 3(1):11:1–11:28, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2011.
3.1/jtse.2011.3.1.1095/jtse.2011.3.1.1095.xml.

Christensen:2011:DCD

[CHP11] Timothy Christensen, Stan Hurn, and Adrian Pagan. Detecting common dynamics in transitory components. Journal of Time Series Econometrics, 3(1):3:1–3:28, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2011.
3.1/jtse.2011.3.1.1088/jtse.2011.3.1.1088.xml.

Cardinali:2010:CLS

[CN10] Alessandro Cardinali and Guy P. Nason. Costationarity of locally stationary time series. Journal of Time Series Econometrics, 2(2): 1:1–1:33, December 2010. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2011.2.2/jtse.2011.2.2.1074/jtse.2011.2.2.1074.
xml.

Chen:2020:TVN

[CP20] Jie Chen and Dimitris N. Politis. Time-varying NoVaS versus GARCH: Point prediction, volatility estimation and prediction intervals. Journal of Time Series Econometrics, 12(2):??, ???? 2020. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2019-0044/html.

Carrara:2022:ECV

[CP22] Aniela Fagundes Carrara and Tiago Luiz Pesquero. The export of commodities and the validity of the export-led growth (ELG) hypothesis for the Brazilian economy: an analysis of the commodity boom period. Journal of Time Series Econometrics, 14(1):87–106, January 2022. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/jtse-2020-0034/html.

Demetrescu:2009:PUR

[Dem09] Matei Demetrescu. Panel unit root testing with nonlinear instruments for infinite-order autoregressive processes. Journal of Time Series Econometrics, 1(2):??, December 2009. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2009.
1.2/jtse.2009.1.2.1009/jtse.2009.1.2.1009.xml.

Dimitriou-Fakalou:2022:DWU

[DF22] Chrysoula Dimitriou-Fakalou. On a different way of understanding the edge-effect for the inference of ARMA-type processes (in Z^{d}). Journal of Time Series Econometrics, 14(1):25–50, January 2022. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2020-0012/html.

Dahl:2011:MVR

[DI11] Christian M. Dahl and Emma Iglesias. Modeling the volatility– return trade-off when volatility may be nonstationary. Journal of Time Series Econometrics, 3(1):10:1–10:32, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2011.
3.1/jtse.2011.3.1.1093/jtse.2011.3.1.1093.xml.

Demos:2019:FST

[DK19] Antonis Demos and Dimitra Kyriakopoulou. Finite-sample theory and bias correction of maximum likelihood estimators in the EGARCH model. Journal of Time Series Econometrics, 11(1):??, January 2019. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL http://www.degruyter.com/view/j/jtse.
2019.11.issue-1/jtse-2018-0010/jtse-2018-0010.xml.

Doojav:2023:FIM

[DL23] Gan-Ochir Doojav and Davaajargal Luvsannyam. Forecasting inflation in Mongolia: a dynamic model averaging approach. Journal of Time Series Econometrics, 15(1):27–48, January 2023. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2020-0021/html.

delBarrioCastro:2011:NTP

[dO11] Tomás del Barrio Castro and Denise R. Osborn. Nonparametric tests for periodic integration. Journal of Time Series Econometrics, 3(1):4:1–4:35, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2011.3.1/jtse.2011.3.1.1090/jtse.2011.
3.1.1090.xml.

Davidson:2015:TLM

[DR15] James Davidson and Dooruj Rambaccussing. A test of the long memory hypothesis based on self-similarity. Journal of Time Series Econometrics, 7(2):115–141, July 2015. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2015.7.issue-
2/jtse-2013-0036/jtse-2013-0036.xml.

Everaert:2010:EIT

[Eve10] Gerdie Everaert. Estimation and inference in time series with omitted I(1) variables. Journal of Time Series Econometrics, 2(2): 2:1–2:26, December 2010. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2011.2.2/jtse.2011.2.2.1054/jtse.2011.2.2.1054.
xml.

Gabrys:2013:MIV

[GHK13] Robertas Gabrys, Siegfried Hörmann, and Piotr Kokoszka. Monitoring the intraday volatility pattern. Journal of Time Series Econometrics, 5(2):87–116, November 2013. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.

degruyter.com/view/j/jtse.2013.5.issue-2/jtse-2012-0006/
jtse-2012-0006.xml.

Gourieroux:2023:TLM

[GJ23] Christian Gourieroux and Joann Jasiak. Temporally local maximum likelihood with application to SIS model. Journal of Time Series Econometrics, 15(2):151–198, July 2023. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2022-0016/html.

Grassi:2010:VUI

[GP10] Stefano Grassi and Tommaso Proietti. Has the volatility of U.S. inflation changed and how? Journal of Time Series Econometrics, 2(1):??, January 2010. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2010.2.1/jtse.2010.2.1.1050/jtse.2010.2.1.1050.
xml.

Golosov:2014:MSR

[GS14] Edward Golosov and Stephen Satchell. Modeling style rotation: Switching and re-switching. Journal of Time Series Econometrics, 6(2):103–128, July 2014. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2014.6.issue-2/jtse-2012-0028/jtse-2012-0028.xml.

Game:2013:CRB

[GW13] Aaron Game and Jason Wu. A covariate residual-based cointegration test applied to the CDS-bond basis. Journal of Time Series Econometrics, 5(2):163–192, November 2013. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2013.5.issue-
2/jtse-2012-0020/jtse-2012-0020.xml.

Zerom:2020:PAP

[GZ20] Jan G. De Gooijer and Dawit Zerom. Penalized averaging of parametric and non-parametric quantile forecasts. Journal of Time Series Econometrics, 12(1):??, January 2020. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL https://www.degruyter.com/view/journals/jtse/12/1/
article-20190021.xml.

Hualde:2012:FSE

[HI12] Javier Hualde and Fabrizio Iacone. First stage estimation of fractional cointegration. Journal of Time Series Econometrics, 4(1): 2:1–2:32, May 2012. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL http://www.degruyter.com/view/j/
jtse.2012.4.issue-1/1941-1928.1129/1941-1928.1129.xml.

Hecq:2016:URB

[HLP16] Alain Hecq, Sébastien Laurent, and Franz C. Palm. On the univariate representation of BEKK models with common factors. Journal of Time Series Econometrics, 8(2):91–113, July 2016. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2016.8.issue-
2/jtse-2015-0002/jtse-2015-0002.xml.

Hendry:2011:EMT

[HM11] David F. Hendry and Grayham E. Mizon. Econometric modelling of time series with outlying observations. Journal of Time Series Econometrics, 3(1):6:1–6:26, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.

degruyter.com/view/j/jtse.2011.3.1/jtse.2011.3.1.1100/
jtse.2011.3.1.1100.xml.

Haldrup:2011:DAO

[HMS11] Niels Haldrup, Antonio Montañes, and Andreu Sansó. Detection of additive outliers in seasonal time series. Journal of Time Series Econometrics, 3(2):2:1–2:20, April 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2011.
3.2/jtse.2011.3.2.1043/jtse.2011.3.2.1043.xml.

Hillebrand:2013:ATR

[HMX13] Eric Hillebrand, Marcelo C. Medeiros, and Junyue Xu. Asymptotic theory for regressions with smoothly changing parameters. Journal of Time Series Econometrics, 5(2):133–162, November 2013. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2013.5.issue-
2/jtse-2012-0024/jtse-2012-0024.xml.

Hunt:2021:GFD

[HPW21] Richard Hunt, Shelton Peiris, and Neville Weber. A general frequency domain estimation method for Gegenbauer processes. Journal of Time Series Econometrics, 13(2):119–144, July 2021. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2019-0031/html.

Hassler:2013:ABT

[HT13] Uwe Hassler and Henghsiu Tsai. Asymptotic behavior of temporal aggregates in the frequency domain. Journal of Time Series Econometrics, 5(1):47–60, May 2013. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2013.5.issue-
1/jtse-2012-0029/jtse-2012-0029.xml.

Iacone:2017:TCM

[ILT17] Fabrizio Iacone, Stephen J. Leybourne, and A. M. Robert Taylor. Testing for a change in mean under fractional integration. Journal of Time Series Econometrics, 9(1):20150006:1–20150006:8, January 2017. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2017.
9.issue-1/jtse-2015-0006/jtse-2015-0006.xml.

Jensen:2010:CQM

[JL10] Anders Tolver Jensen and Theis Lange. On convergence of the QMLE for misspecified GARCH models. Journal of Time Series Econometrics, 2(1):??, January 2010. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.

degruyter.com/view/j/jtse.2010.2.1/jtse.2010.2.1.1034/
jtse.2010.2.1.1034.xml.

Jansson:2011:NEL

[JN11] Michael Jansson and Morten Ørregaard Nielsen. Nearly efficient likelihood ratio tests for seasonal unit roots. Journal of Time Series Econometrics, 3(1):5:1–5:21, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2011.
3.1/jtse.2011.3.1.1096/jtse.2011.3.1.1096.xml.

Javed:2017:TBD

[JP17] Farrukh Javed and Krzysztof Podgórski. Tail behavior and dependence structure in the APARCH model. Journal of Time Series Econometrics, 9(2):20160002:1–20160002:48, July 2017. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2017.9.issue-
2/jtse-2016-0002/jtse-2016-0002.xml.

Johansen:2011:GTE

[JS11] Søren Johansen and Anders R. Swensen. On a graphical technique for evaluating some rational expectations models. Journal of Time Series Econometrics, 3(1):9:1–9:29, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2011.
3.1/jtse.2011.3.1.1089/jtse.2011.3.1.1089.xml.

Jiang:2018:VML

[JX18] Zhengjun Jiang and Weixuan Xia. Volatility modeling with leverage effect under Laplace errors. Journal of Time Series Econometrics, 10(1):20160019:1–20160019:28, January 2018. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2018.
10.issue-1/jtse-2016-0019/jtse-2016-0019.xml.

Kawakatsu:2023:SFR

[Kaw23] Hiroyuki Kawakatsu. Simple factor realized stochastic volatility models. Journal of Time Series Econometrics, 15(1):79–110, January 2023. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/jtse-2021-0049/html.

Kokoszka:2019:RAC

[KMSZ19] Piotr Kokoszka, Hong Miao, Stilian Stoev, and Ben Zheng. Risk analysis of cumulative intraday return curves. Journal of Time Series Econometrics, 11(2):??, July 2019. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL

https://www.degruyter.com/view/j/jtse.2019.11.issue-2/
jtse-2018-0011/jtse-2018-0011.xml.

Kock:2011:FUA

[Koc11] Anders Bredahl Kock. Forecasting with universal approximators and a learning algorithm. Journal of Time Series Econometrics, 3 (3):3:1–3:32, October 2011. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2011.3.issue-3/1941-1928.1084/1941-1928.1084.xml.

Kolios:2019:PBC

[Kol19] Bill Kolios. Political business cycles in Australia elections and party ideology. Journal of Time Series Econometrics, 11(2):??, July 2019. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL https://www.degruyter.com/view/j/jtse.
2019.11.issue-2/jtse-2017-0012/jtse-2017-0012.xml.

Korhonen:2021:BDR

[KP21] Marko Korhonen and Mikko Puhakka. The behavior of divorce rates: A smooth transition regression approach. Journal of Time Series Econometrics, 13(1):1–19, January 22, 2021. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2019-0018/html.

Kristensen:2009:AQN

[KR09] Dennis Kristensen and Anders Rahbek. Asymptotics of the QMLE for non-linear ARCH models. Journal of Time Series Econometrics, 1(1):??, January 2009. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2009.1.1/jtse.2009.1.1.1001/jtse.2009.1.1.1001.
xml.

Knight:2011:SNR

[KS11] John Knight and Stephen Satchell. Some new results for threshold AR(1) models. Journal of Time Series Econometrics, 3(2):1:1– 1:42, April 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL http://www.degruyter.com/view/j/jtse.
2011.3.2/jtse.2011.3.2.1085/jtse.2011.3.2.1085.xml.

Khan:2017:AFB

[KSJ17] Naushad Mamode Khan, Yuvraj Sunecher, and Vandna Jowaheer. Analyzing the full BINMA time series process using a robust GQL approach. Journal of Time Series Econometrics, 9(2):20150019:1– 20150019:12, July 2017. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2017.9.issue-2/jtse-2015-0019/jtse-2015-0019.xml.

Khan:2018:FOD

[KSJ18] Naushad Mamode Khan, Yuvraj Sunecher, and Vandna Jowaheer. A flexible observation-driven stationary bivariate negative binomial INAR(1) with non-homogeneous levels of overdispersion. Journal of Time Series Econometrics, 10(2): 20160028:1–20160028:8, July 2018. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2018.10.issue-2/jtse-2016-0028/jtse-2016-
0028.xml.

Kurozumi:2015:TMS

[Kur15] Eiji Kurozumi. Testing for multiple structural changes with nonhomogeneous regressors. Journal of Time Series Econometrics, 7 (1):1–35, January 2015. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2015.7.issue-1/jtse-2012-0019/jtse-2012-0019.xml.

Larsson:2015:HCF

[Lar15] Rolf Larsson. How close is a fractional process to a random walk with drift? Journal of Time Series Econometrics, 7(2):217–234, July 2015. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2015.
7.issue-2/jtse-2013-0032/jtse-2013-0032.xml.

Laurini:2013:HDC

[Lau13] Márcio Poletti Laurini. A hybrid data cloning maximum likelihood estimator for stochastic volatility models. Journal of Time Series Econometrics, 5(2):193–229, November 2013. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2013.5.issue-
2/jtse-2012-0025/jtse-2012-0025.xml.

Livio:2020:IPI

[LBN20] Tito Lívio, Marcelo Bourguignon, and Fernando Nascimento. INAR(1) processes with inflated-parameter generalized power series innovations. Journal of Time Series Econometrics, 12(2):??, ???? 2020. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/jtse-2019-0033/html.

Liu-Evans:2023:IEP

[LE23] Gareth Liu-Evans. Improving the estimation and predictions of small time series models. Journal of Time Series Econometrics, 15(1):1–26, January 2023. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://www.degruyter.

com/document/doi/10.1515/jtse-2021-0051/html.

Lee:2019:NNM

[Lee19] Jinu Lee. A neural network method for nonlinear time series analysis. Journal of Time Series Econometrics, 11(1):??, January 2019. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2019.
11.issue-1/jtse-2016-0011/jtse-2016-0011.xml.

Liu-Evans:2012:BJC

[LEP12] Gareth D. Liu-Evans and Garry D. A. Phillips. Bootstrap, jackknife and COLS: Bias and mean squared error in estimation of autoregressive models. Journal of Time Series Econometrics, 4(2):1:1–1:33, November 2012. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2012.4.issue-
2/1941-1928.1122/1941-1928.1122.xml.

Lips:2017:DTS

[Lip17] Johannes Lips. Do they still matter? — Impact of fossil fuels on electricity prices in the light of increased renewable generation. Journal of Time Series Econometrics, 9(2):20160018:1– 20160018:30, July 2017. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2017.9.issue-2/jtse-2016-0018/jtse-2016-0018.xml.

Li:2010:NSL

[LL10] Dong Li and Canh Le. Nonlinearity and spatial lag dependence: Tests based on double-length regressions. Journal of Time Series Econometrics, 2(1):??, January 2010. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.

degruyter.com/view/j/jtse.2010.2.1/jtse.2010.2.1.1039/
jtse.2010.2.1.1039.xml.

Lanne:2011:NAE

[LS11] Markku Lanne and Pentti Saikkonen. Noncausal autoregressions for economic time series. Journal of Time Series Econometrics, 3 (3):2:1–2:32, October 2011. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2011.3.issue-3/1941-1928.1080/1941-1928.1080.xml.

Lima:2010:TUR

[LX10] Luiz Renato Lima and Zhijie Xiao. Testing unit root based on partially adaptive estimation. Journal of Time Series Econometrics, 2(1):??, January 2010. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2010.2.1/jtse.2010.2.1.1038/jtse.2010.2.1.1038.
xml.

Luetkepohl:2011:FAI

[LX11] Helmut Luetkepohl and Fang Xu. Forecasting annual inflation with seasonal monthly data: Using levels versus logs of the underlying price index. Journal of Time Series Econometrics, 3(1): 7:1–7:23, January 2011. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2011.3.1/jtse.2011.3.1.1094/jtse.2011.3.1.1094.
xml.

Lee:2013:TNN

[LXZ13] Tae-Hwy Lee, Zhou Xi, and Ru Zhang. Testing for neglected nonlinearity using artificial neural networks with many randomized hidden unit activations. Journal of Time Series Econometrics, 5 (1):61–86, May 2013. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL http://www.degruyter.com/view/j/
jtse.2013.5.issue-1/jtse-2012-0021/jtse-2012-0021.xml.

Mainassara:2022:ESM

[MA22a] Yacouba Boubacar Maïnassara and Abdoulkarim Ilmi Amir. Estimating SPARMA models with dependent error terms. Journal of Time Series Econometrics, 14(2):141–174, July 2022. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2021-0022/html.

Mainassara:2022:GFT

[MA22b] Yacouba Boubacar Maïnassara and Abdoulkarim Ilmi Amir. Goodness-of-fit tests for SPARMA models with dependent error terms. Journal of Time Series Econometrics, 14(2):107–140, July 2022. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2022-0002/html.

Man:2010:EFG

[Man10] Kasing Man. Extended fractional Gaussian noise and simple ARFIMA approximations. Journal of Time Series Econometrics, 2(1):??, January 2010. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2010.2.1/jtse.2010.2.1.1063/jtse.2010.2.1.1063.
xml.

Mansur:2024:CPI

[Man24] Alfan Mansur. Commodity price and Indonesian fiscal policy: an SVAR analysis with non-Gaussian errors. Journal of Time Series Econometrics, 16(1):29–66, January 2024.

CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2023-0037/html.

Marfatia:2021:MHP

[Mar21] Hardik A. Marfatia. Modeling house price synchronization across the U.S. states and their time-varying macroeconomic linkages. Journal of Time Series Econometrics, 13(1):73–117, January 22, 2021. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2017-0014/html.

Miller:2010:NIL

[Mil10] J. Isaac Miller. A nonlinear IV likelihood-based rank test for multivariate time series and long panels. Journal of Time Series Econometrics, 2(1):??, January 2010. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2010.2.1/jtse.2010.2.1.1057/jtse.2010.
2.1.1057.xml.

Mallory:2012:TCP

[ML12] Mindy Mallory and Sergio H. Lence. Testing for cointegration in the presence of moving average errors. Journal of Time Series Econometrics, 4(2):2:1–2:66, November 2012. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2012.4.issue-
2/1941-1928.1124/1941-1928.1124.xml.

McElroy:2014:OSE

[MM14] Tucker S. McElroy and Agustin Maravall. Optimal signal extraction with correlated components. Journal of Time Series Econometrics, 6(2):237–273, July 2014. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2014.6.issue-2/jtse-2013-0016/jtse-2013-
0016.xml.

Montes-Rojas:2022:EIR

[MR22] Gabriel Montes-Rojas. Estimating impulse-response functions for macroeconomic models using directional quantiles. Journal of Time Series Econometrics, 14(2):199–225, July 2022. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2021-0002/html.

Morettin:2011:WEC

[MTCdM11] Pedro A. Morettin, Clelia M. C. Toloi, Chang Chiann, and José C. S. de Miranda. Wavelet estimation of copulas for time series. Journal of Time Series Econometrics, 3(3):4:1–4:31, October 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2011.
3.issue-3/1941-1928.1033/1941-1928.1033.xml.

McElroy:2010:SER

[MW10] Tucker McElroy and Marc Wildi. Signal extraction revision variances as a goodness-of-fit measure. Journal of Time Series Econometrics, 2(1):??, January 2010. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2010.2.1/jtse.2010.2.1.1012/jtse.2010.
2.1.1012.xml.

Milunovich:2013:ISV

[MY13] George Milunovich and Minxian Yang. On identifying structural VAR models via ARCH effects. Journal of Time Series Econometrics, 5(2):117–131, November 2013. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2013.5.issue-
2/jtse-2013-0010/jtse-2013-0010.xml.

Ng:2009:SIV

[NB09] Serena Ng and Jushan Bai. Selecting instrumental variables in a data rich environment. Journal of Time Series Econometrics, 1(1):??, January 2009. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2009.1.1/jtse.2009.1.1.1014/jtse.2009.1.1.1014.
xml.

Nguimkeu:2016:IST

[Ngu16] Pierre Nguimkeu. An improved selection test between autoregressive and moving average disturbances in regression models. Journal of Time Series Econometrics, 8(1):41–54, January 2016. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2016.8.issue-
1/jtse-2014-0036/jtse-2014-0036.xml.

Nonejad:2016:PMC

[Non16] Nima Nonejad. Particle Markov chain Monte Carlo techniques of unobserved component time series models using Ox. Journal of Time Series Econometrics, 8(1):55–90, January 2016. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2016.8.issue-
1/jtse-2013-0024/jtse-2013-0024.xml.

Olivares:2021:ECD

[OG21] Daniel González Olivares and Isai Guizar. Estimation of continuous and discrete time co-integrated systems with stock and flow variables. Journal of Time Series Econometrics, 13(2):145–186, July 2021. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/jtse-2019-0026/html.

Okui:2014:AUE

[Oku14] Ryo Okui. Asymptotically unbiased estimation of autocovariances and autocorrelations with panel data in the presence of individual and time effects. Journal of Time Series Econometrics, 6(2):129– 181, July 2014. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL http://www.degruyter.com/view/j/jtse.
2014.6.issue-2/jtse-2013-0017/jtse-2013-0017.xml.

Otunuga:2019:LLA

[OLL19] Olusegun M. Otunuga, Gangaram S. Ladde, and Nathan G. Ladde. Local lagged adapted generalized method of moments: An innovative estimation and forecasting approach and its applications. Journal of Time Series Econometrics, 11(1):??, January 2019. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2019.
11.issue-1/jtse-2016-0024/jtse-2016-0024.xml.

Ollech:2021:SAD

[Oll21] Daniel Ollech. Seasonal adjustment of daily time series. Journal of Time Series Econometrics, 13(2):235–264, July 2021. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2020-0028/html.

Pauwels:2012:TSC

[PCG12] Laurent L. Pauwels, Felix Chan, and Tommaso Mancini Griffoli. Testing for structural change in heterogeneous panels with an application to the euro’s trade effect. Journal of Time Series Econometrics, 4(2):3:1–3:33, November 2012. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2012.4.issue-
2/1941-1928.1141/1941-1928.1141.xml.

Porto:2012:RAE

[PMA12] Rogério F. Porto, Pedro A. Morettin, and Elisete C. Q. Aubin. Regression with autocorrelated errors using design-adapted Haar wavelets. Journal of Time Series Econometrics, 4(1):4:1–4:30, May 2012. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2012.
4.issue-1/1941-1928.1067/1941-1928.1067.xml.

Poissonnier:2018:CLM

[Poi18] Aurélien Poissonnier. The Chow-Lin method extended to dynamic models with autocorrelated residuals. Journal of Time Series Econometrics, 10(1):20160007:1–20160007:16, January 2018. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2018.
10.issue-1/jtse-2016-0007/jtse-2016-0007.xml.

Pollock:2009:SFA

[Pol09] Stephen D. S. G. Pollock. Statistical Fourier analysis: Clarifications and interpretations. Journal of Time Series Econometrics, 1(1):??, January 2009. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2009.1.1/jtse.2009.1.1.1004/jtse.2009.1.1.1004.
xml.

Pollock:2014:CSS

[Pol14] D. S. G. Pollock. Cycles, syllogisms and semantics: Examining the idea of spurious cycles. Journal of Time Series Econometrics, 6(1):81–102, January 2014. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2014.6.issue-1/jtse-2012-0033/jtse-2012-0033.xml.

Parker:2015:TBB

[PPP15] Cameron C. Parker, Efstathios Paparoditis, and Dimitris Politis. Tapered block bootstrap for unit root testing. Journal of Time Series Econometrics, 7(1):37–67, January 2015. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2015.7.issue-
1/jtse-2013-0033/jtse-2013-0033.xml.

Perron:2011:IIT

[PR11] Pierre Perron and Linxia Ren. On the irrelevance of impossibility theorems: The case of the long-run variance. Journal of Time Series Econometrics, 3(3):1:1–1:34, October 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2011.3.issue-
3/1941-1928.1062/1941-1928.1062.xml.

Quineche:2021:CAW

[Qui21] Ricardo Quineche. Consumption, aggregate wealth and expected stock returns: An FCVAR approach. Journal of Time Series Econometrics, 13(1):21–42, January 22, 2021. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://

www.degruyter.com/document/doi/10.1515/jtse-2020-0029/
html.

Reed:2010:PEG

[RW10] W. Robert Reed and Rachel Webb. The PCSE estimator is good — just not as good as you think. Journal of Time Series Econometrics, 2(1):??, January 2010. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.degruyter.

com/view/j/jtse.2010.2.1/jtse.2010.2.1.1032/jtse.2010.
2.1.1032.xml.

Sanhaji:2017:TNC

[San17] Bilel Sanhaji. Testing for nonlinearity in conditional covariances. Journal of Time Series Econometrics, 9(2):20160010:1– 20160010:22, July 2017. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2017.9.issue-2/jtse-2016-0010/jtse-2016-0010.xml.

Shang:2020:CHE

[Sha20] Han Lin Shang. A comparison of Hurst exponent estimators in long-range dependent curve time series. Journal of Time Series Econometrics, 12(1):??, January 2020. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL https://www.degruyter.com/view/journals/jtse/12/1/
article-20190009.xml.

Simos:2012:EDC

[Sim12] Theodore Simos. On the exact discretization of a continuous time AR(1) model driven by either long memory or antipersistent innovations: A fractional algebra approach. Journal of Time Series Econometrics, 4(2):5:1–5:24, November 2012. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2012.4.issue-
2/1941-1928.1145/1941-1928.1145.xml.

Singh:2016:IMC

[Sin16] Tarlok Singh. International mobility of capital in the United States: robust evidence from time-series tests. Journal of Time Series Econometrics, 8(2):193–249, July 2016. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2016.8.issue-
2/jtse-2014-0005/jtse-2014-0005.xml.

Sunecher:2019:MDB

[SKJ19] Yuvraj Sunecher, Naushad Mamode Khan, and Vandna Jowaheer. Modelling with dispersed bivariate moving average processes. Journal of Time Series Econometrics, 11(1):??, January 2019. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2019.
11.issue-1/jtse-2018-0009/jtse-2018-0009.xml.

Skrobotov:2014:BCK

[Skr14] Anton Skrobotov. Bias correction of KPSS test with structural break for reducing of size distortion. Journal of Time Series Econometrics, 6(1):33–61, January 2014. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http://www.

degruyter.com/view/j/jtse.2014.6.issue-1/jtse-2012-0031/
jtse-2012-0031.xml.

Skrobotov:2018:TBI

[Skr18] Anton Skrobotov. On trend breaks and initial condition in unit root testing. Journal of Time Series Econometrics, 10(1):20160014:1–20160014:14, January 2018. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http:

//www.degruyter.com/view/j/jtse.2018.10.issue-1/jtse-
2016-0014/jtse-2016-0014.xml.

Symeonides:2017:SCS

[SKT17] Spyridon D. Symeonides, Yiannis Karavias, and Elias Tzavalis. Size corrected significance tests in seemingly unrelated regressions with autocorrelated errors. Journal of Time Series Econometrics, 9(1):20150014:1–20150014:14, January 2017. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2017.9.issue-
1/jtse-2015-0014/jtse-2015-0014.xml.

Smith:2012:MBR

[Smi12] Aaron Smith. Markov breaks in regression models. Journal of Time Series Econometrics, 4(1):3:1–3:35, May 2012. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2012.4.issue-
1/1941-1928.1111/1941-1928.1111.xml.

Sancetta:2009:FPV

[SN09] Alessio Sancetta and Arina Nikandrova. Forecasting and prequential validation for time varying meta-elliptical distributions. Journal of Time Series Econometrics, 1(2):??, December 2009. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2009.
1.2/jtse.2009.1.2.1005/jtse.2009.1.2.1005.xml.

Sollis:2016:FRR

[Sol16] Robert Sollis. Fixed and recursive right-tailed Dickey-Fuller tests in the presence of a break under the null. Journal of Time Series Econometrics, 8(1):1–19, January 2016. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2016.8.issue-
1/jtse-2013-0004/jtse-2013-0004.xml.

Tayanagi:2023:FAD

[TK23] Toshikazu Tayanagi and Eiji Kurozumi. In-fill asymptotic distribution of the change point estimator when estimating breaks one at a time. Journal of Time Series Econometrics, 15(2):111–149, July 2023. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL https://www.degruyter.com/document/doi/
10.1515/jtse-2022-0013/html.

Tripathi:2021:ERF

[TKI21] Manas Tripathi, Saurabh Kumar, and Sarveshwar Kumar Inani. Exchange rate forecasting using ensemble modeling for better policy implications. Journal of Time Series Econometrics, 13(1):43– 71, January 22, 2021. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL https://www.degruyter.com/document/
doi/10.1515/jtse-2020-0013/html.

Trimbur:2017:SEN

[TM17] Thomas Trimbur and Tucker McElroy. Signal extraction for nonstationary time series with diverse sampling rules. Journal of Time Series Econometrics, 9(1):20140026:1–20140026:37, January 2017. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2017.9.issue-
1/jtse-2014-0026/jtse-2014-0026.xml.

Taufemback:2022:RTM

[TTS22] Cleiton G. Taufemback, Victor Troster, and Muhammad Shahbaz. A robust test for monotonicity in asset returns. Journal of Time Series Econometrics, 14(1):1–24, January 2022. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2019-0068/html.

Tofoli:2019:DVC

[TZCP19] Paula V. Tófoli, Flávio A. Ziegelmann, Osvaldo Candido, and Pedro L. Valls Pereira. Dynamic D-vine copula model with applications to value-at-risk (VaR). Journal of Time Series Econometrics, 11(2):??, July 2019. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL https://www.degruyter.com/view/
j/jtse.2019.11.issue-2/jtse-2017-0016/jtse-2017-0016.
xml.

Vafiadis:2015:FVR

[Vaf15] Nikolaos Vafiadis. Forecasting volatility and the risk-return tradeoff: an application on the Fama-French benchmark market return. Journal of Time Series Econometrics, 7(2):181–216, July 2015. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2015.7.issue-
2/jtse-2012-0018/jtse-2012-0018.xml.

Ventosa-Santaularia:2010:TDT

[VSGZ10] Daniel Ventosa-Santaulària and Manuel Gómez-Zaldívar. Testing for a deterministic trend when there is evidence of unit root. Journal of Time Series Econometrics, 2(2):3:1–3:24, December 2010. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL http://www.degruyter.com/view/j/jtse.2011.
2.2/jtse.2011.2.2.1013/jtse.2011.2.2.1013.xml.

Wang:2014:BPO

[Wan14] Liqiong Wang. Bootstrap point optimal unit root tests. Journal of Time Series Econometrics, 6(1):1, January 2014. CO-DEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic).

URL http://www.degruyter.com/view/j/jtse.2014.6.issue-
1/jtse-2013-0006/jtse-2013-0006.xml.

White:2011:CTT

[WG11] Halbert White and Clive W. J. Granger. Consideration of trends in time series. Journal of Time Series Econometrics, 3(1):2:1–2:40, January 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928

(electronic). URL http://www.degruyter.com/view/j/jtse.
2011.3.1/jtse.2011.3.1.1092/jtse.2011.3.1.1092.xml.

Wang:2011:EAP

[WH11] Shin-Huei Wang and Christian Hafner. Estimating autocorrelations in the presence of deterministic trends. Journal of Time Series Econometrics, 3(2):4:1–4:25, April 2011. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL http:/

/www.degruyter.com/view/j/jtse.2011.3.2/jtse.2011.3.2.
1022/jtse.2011.3.2.1022.xml.

Wang:2013:RTM

[WH13] Cindy Shin-Huei Wang and Cheng Hsiao. Real-time monitoring test for realized volatility. Journal of Time Series Econometrics, 5 (1):1–24, May 2013. CODEN ???? ISSN 2194-6507 (print), 1941-

1928 (electronic). URL http://www.degruyter.com/view/j/
jtse.2013.5.issue-1/jtse-2012-0014/jtse-2012-0014.xml.

Wildi:2016:ORT

[WM16] Marc Wildi and Tucker McElroy. Optimal real-time filters for linear prediction problems. Journal of Time Series Econometrics, 8 (2):155–192, July 2016. CODEN ???? ISSN 2194-6507 (print),

1941-1928 (electronic). URL http://www.degruyter.com/view/
j/jtse.2016.8.issue-2/jtse-2014-0019/jtse-2014-0019.xml.

Weiss:2020:CMA

[WSAF20] Christian Weiß, Lukas Scherer, Boris Aleksandrov, and Martin Feld. Checking model adequacy for count time series by using Pearson residuals. Journal of Time Series Econometrics, 12(1):??, January 2020. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (electronic). URL https://www.degruyter.

com/view/journals/jtse/12/1/article-20180018.xml.

Xu:2024:QML

[Xu24] Yongdeng Xu. Quasi maximum likelihood estimation of vector multiplicative error model using the ECCC-GARCH representation. Journal of Time Series Econometrics, 16(1):1–27, January 2024. CODEN ???? ISSN 2194-6507 (print), 1941-1928 (elec-

tronic). URL https://www.degruyter.com/document/doi/10.
1515/jtse-2022-0018/html.